OpenStudy (anonymous):

  Which two expressions are equivalent? A. 4.3(14x+2.5) B. 14(4.3x+2.5) C. 20.6x+28+39.6x+7 D. 60.2x+2.5 Medal and fan for best answer. I am not good at figuring out equivalent expressions.

OpenStudy (mathstudent55):

You need to use the distributive property in each choice A and B. Combine like terms in choice C. Then compare all results including choice D.

OpenStudy (mathstudent55):

Recall the distributive property: a(b + c) = ab + ac Example: 5(2x + 4) = 5 * 2x + 5 * 4 = 10x + 20

OpenStudy (mathstudent55):

Can you do the distributive property on choice A? What do you get? A. 4.3(14x+2.5)

OpenStudy (anonymous):

Ok so I have 4.3*2.5, but i'm having trouble multiplying 4.3 *14x. I can't seem to figure out how to multiply with the variable number multiplication thingy.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@carolinar7

OpenStudy (anonymous):

@sweetburger @Elsa213 @YanaSidlinskiy @misssunshinexxoxo

OpenStudy (anonymous):

Anyone? I would like help plz. My schoolwork is taken me all day and I wanna b done with it.

jimthompson5910 (jim_thompson5910):

to multiply 4.3 *14x, you drop the x for now and just evaluate 4.3 *14 you then tack on x to the answer

jimthompson5910 (jim_thompson5910):

example: 2*8x = 16x because 2*8 = 16

OpenStudy (anonymous):

ty!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (mathstudent55):

Just use a calculator for the multiplications with decimals.

OpenStudy (mathstudent55):

A. 4.3(14x+2.5) = 4.3 * 14x + 4.3 * 2.5 Use a calculator for 4.3 * 14 and for 4.3 * 2.5

OpenStudy (mathstudent55):

A. 4.3(14x+2.5) = 4.3 * 14x + 4.3 * 2.5 = 60.2x + 10.75

OpenStudy (mathstudent55):

You can tell already that A and D are not the same.

OpenStudy (mathstudent55):

Now let's try B. We need to distribute again. B. 14(4.3x+2.5) = 14 * 4.3x + 14 * 2.5 Can you finish the expression above by doing the two multiplications?

OpenStudy (anonymous):

I got 95.2 by adding the sums of the two multiplications.

OpenStudy (mathstudent55):

No. You can add the two quantities. One has x and one has no x. The are not like quantities, so you can add them. B. 14(4.3x+2.5) = 14 * 4.3x + 14 * 2.5 = 60.2x + 35

OpenStudy (mathstudent55):

So far we have: A. 60.2x + 10.75 B. 60.2x + 35 C. D. 60.2x + 2.5 None of them are equivalent. We must find what C is equal to.

OpenStudy (mathstudent55):

C. 20.6x+28+39.6x+7 You can only combine like terms. Terms with x are like terms. You can combine them together. Terms with no x are like terms. You can combine them together. You cannot combine together terms with x and terms without x.

OpenStudy (mathstudent55):

Examples: 2x + 5x = 7x 2 + 5x = 2 + 5x 5 + 2x + 3 + 4x = 5 + 3 + 2x + 4x = 8 + 6x

OpenStudy (mathstudent55):

Do you understand the idea of like and unlike terms now?

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

Let's try C. now. First we move the like terms together. Then we combine like terms. C. 20.6x+28+39.6x+7 = 20.6x + 39.6x + 28 + 7 Ok so far? All I did was rearrange the terms.

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

Now we combine the like terms with x: 20.6x + 39.6x = 60.2x and we combine the like terms without x: 28 + 7 = 35 We get: C. 20.6x+28+39.6x+7 = 20.6x + 39.6x + 28 + 7 = 60.2x + 35

OpenStudy (mathstudent55):

Now we can write what all the choices are equal to: A. 60.2x + 10.75 B. 60.2x + 35 C. 60.2x + 35 D. 60.2x + 2.5 Now it's obvious which 2 expressions are equivalent.

OpenStudy (anonymous):

Ty for your help and sry I went in and out of the conversation. There is a weird glitch that wont let me type after I send a reply. Thank you!

OpenStudy (mathstudent55):

You're welcome.